# Deep clustering using adversarial net based clustering loss

### v2

@@ IV-A Training the discriminator
$\begin{array}[]{c}L_{2}=\left\{\frac{1}{N}\sum_{n=1}^{N}\;\underset{x^{\left(n\right)}\sim p_{data}}{E}\left[\ln\left(1-D\left(G(x^{\left(n\right)})\right)\right)\right]\right\}\end{array}$ (7)
Neural Network B: The deep clustering parameters $\theta=\{\eta,\tau,\zeta\}$ are defined as
$\begin{array}[]{c}\eta_{k}=\frac{1}{N}\sum_{n=1}^{N}\;z^{(n)}\in k\\ \tau_{k}=\left(\frac{1}{N}\sum_{n=1}^{N}\;\left(z^{(n)}-\eta_{k}\right)^{2}\right)^{-1}\in k\\ \zeta^{(n)}=\arg\max_{k}\left\{\;\ln p(z^{(n)}\mid\eta_{k},\tau_{k})\;\right\}\forall k\end{array}$ (8)
Since the samples in the latent space displaces each time the encoder weight changes, GMM has to be re-computed each time the weight is updated. After sampling from GMM i.e. $z\sim p(z\mid\theta^{*})$ and setting $T=1$ to $D$, we perform weight update using eqn (9).
$\begin{array}[]{c}L_{1}=\left\{\frac{1}{N}\sum_{n=1}^{N}\;\underset{z^{\left(n\right)}\sim p(z|\theta^{*})}{E}\left[\ln D\left(z^{\left(n\right)}\right)\right]\right\}\end{array}$ (9)
IV-B Training the encoder
The encoder $q(z\mid x)$ maps $x$ to $z$ in the latent space. However, we desire to minimize between $q(z\mid x)$ and $p(z\mid\theta^{*}).$ Meaning that we desire to produce a sample $z$ from $q(z\mid x)$ that is as close as possible to the sampling from $p(z\mid\theta^{*})$.
Neural Network R: There is only the discriminator path which updates the encoder weight in Fig 1. Given $N$ samples, eqn (10) computes the SGA weight update for the encoder for $T=0$.
$\begin{array}[]{c}L_{2}=\left\{\frac{1}{N}\sum_{n=1}^{N}\;\underset{x^{\left(n\right)}\sim p_{data}}{E}\left[\ln\left(1-D\left\{G(x^{\left(n\right)})\right\}\right)\right]\right\}\end{array}$ (10)
V Experiments
We selected some benchmarked datasets used by deep clustering for our experiments in Tables II-VI. There are several factors affecting deep clustering. i) The clustering algorithm used in the latent space e.g. Kmeans or GMM. ii) The number of hidden layers and nodes of the neural network. iii) The gradient ascent method used for training the loss functions. iv) The activation functions used.
For DCAN, we use $tanh$ for the hidden layers and $sigmoid$ for the output. We set the clustering iteration to one in Table II and a sufficient statistics of at least 600 points in the raw latent space to estimate the cluster parameters. The number of cluster is the same as the number of classes. We use a minibatch size of $N=16$ samples, 500 iterations and we use gradient ascent with momentum with a learning rate between $10^{-3}$ and $10^{-5}$. We use accuracy (ACC) and normalized mutual information (NMI) [15] to evaluate the performance of our clustering.
There is a tradeoff between clustering accuracy and the dimension reduction representation in the latent space for deep clustering. As compared to most non deep clustering objectives which use a bottleneck latent space (e.g. $Z=10$), we found that having a large dimension for the latent state (e.g. $Z=128$) is crucial for clustering. This is because when the dimensionality is not sufficiently large, there will be alot of overlapping clusterings in the latent space. Empirically, we observed that when the dimension of the latent space is too small, the accuracy obtained will be much lower. This effect can vary between 10 to 20% accuracy for most of the datasets we have tried.
SVHN: A real images dataset containing 10 digits class of house numbers but it is a more complex and challenging dataset than MNIST [16]. The images in SVHN are natural and not clean who have much variance in illumination and distraction. We took a 1K samples per class for 10 classes or 10K random samples in total. We are using Resnet18 [17] as the image feature extraction with a dimension of 512. For this experiment, we use an autoencoder with a single hidden layer with a size of 256 neurons for the encoder (and vice-versa for the decoder) and the latent space has 128 neurons. In Table 2, the average ACC of Kmeans in the original feature space is 0.6815. The latent space trained by reconstruction loss showed clustering performance at an average ACC of 0.7373. The latent space of VAE showed better result when performing Kmeans at 0.7663. But it could not achieve the same performance gained by deep clustering approaches. When performing Kmeans in the latent space of ABC, the average ACC is at 0.8872. Lastly, DCAN performed better result than ABC at 0.9047.
USPS: This is a simple 10 digits class dataset that does not have background noise like SVHN. The training size is 7291 and the testing size is 2007. The raw pixel values of each 16x16 image is flattened to 256 and used as inputs to all the comparisons. Our setup uses $256-196-128$ for the encoder and $128-16-1$ for the discriminator. Kmeans or GMM in the latent space of deep clustering methods such as DC-GMM, ABC, DEC and DCAN outperformed other latent spaces such as using reconstruction loss. This is seen when comparing ABC at 0.715 to AE at 0.6081 in Table III. DCAN outperforms all the methods at $0.8276$.
@@ V Experiments
MNIST: A well cited 10 digit classes dataset with no background but it has more samples than USPS. The train and test sample size is at 50K and 10K. Our settings are $196-384-256$ for the encoder and $256-16-1$ for the decoder. The input dimension is originally at $28\times 28$ but we have downsampled it to $14\times 14$ and flattened to 196. We have also experimented using the original image at $28\times 28$ and flattened to 384 for the encoder at $784-512-384-256$ but the accuracy is slightly poorer. In Table IV, DCAN using raw pixel information obtained $0.8565$ for ACC which is on par or better than the unsupervised deep learning methods including DASC, DEC, DCN and DC-GMM. However on MNIST, most deep clustering methods could not obtain the performance of recent deep learning methods such as KINGDRA, DAC and IMSAT as their goals are much more ACC result oriented. For KINGDRA, it combines several ideas together including ladder network [18] and psuedo-label learning [19] for the neural network and using ensemble clustering. In DAC, their approach is based on pairwise classification. In IMSAT-RPT, the neural network was trained by augmenting the training dataset. In this paper however, we are mainly focused on improving the concept of the Euclidean distance based deep learning using ABC. DCAN penalizes samples for being an outlier from their cluster centers and does so uniquely by using the discriminator to penalize the encoder.
The original objective of deep clustering is specifically aimed at clustering. In some datasets, the latent space of this objective do not necessarily co-serve as the most important factor for raw pixel feature extraction. Thus, we used ResNet18 pretrained model as a feature extractor for MNIST, prior to deep clustering. The new ACC for DCAN improved to 0.995 and in fact outperforms KINGDRA as well as all the recent non deep clustering methods.
CIFAR10: A 10 classes object categorization dataset with 50K training and 10K testing of real images. On CIFAR10, most deep clustering methods have difficulty performing end-to-end learning. As compared to deep ConvNet which utilizes convolutional neural network and supervised learning for a specific task (i.e. feature extraction), deep clustering typically rely on fewer hidden layer in the encoder for end-to-end learning (i.e. both feature extraction and deep clustering). Thus, we followed the experiment setup in [20] by using the “avg pool” layer of ResNet50 as a feature extractor with a dimension of 2048. The ResNet50 [17] model is pretrained on ImageNet. In Table V, DCAN uses $2048-1024-512-128$ and $128-16-1$ respectively for the encoder and discriminator. Despite the stronger performance of DAC, IMSAT and KINGDRA on MNIST, we outperformed these comparisons on CIFAR10 at $0.5844$. This is despite the fact that both KINGDRA and IMSAT also use ResNet50 pretrained on ImageNet.
Reuters-10k: The Reuters-10k dataset according to [21, 22] contains 10K samples with 4 classes and the feature dimension is 2000. End-to-end learning is performed on this dataset. We compared DCAN to other unsupervised deep learning approaches. DCAN uses $2000-100-500$ and $500-100-1$ respectively for the encoder and discriminator. In Table VI, we observed that Kmeans alone can achieve an ACC of 0.6018 on the raw feature dimension. Most deep learning methods can achieve between 0.69 to 0.73 on this dataset. DCAN was able to obtain the best ACC at 0.7867.
INITIALIZATION:
initialize clusters in latent space using Kmeans
initialize encoder and discriminator weights
MAIN:
for epoch=1:EPOCH
1) draw a minibatch $x=\left\{x_{n}\right\}_{n=1}^{N}\in\mathbb{R^{\mathbf{\mathrm{I}}}}$ which contains data space sample
2) forward pass $x$ through encoder to obtain a matrix, $LATENT\in\mathbb{R^{\mathbf{\mathrm{N\times I}}}}$ in the latent space
for iteration=1:ITERATION
3) train Kmeans parameters $\eta,\zeta$ using $LATENT$
end
for minibatch=1:MINIBATCH
4) draw a sample $x_{n}$
5) discriminator: compute network weight using $L_{1}$ and $\eta,\tau,\zeta$ for $T=1$
6) discriminator: compute network weight using $L_{2}$ for $T=0$
7) encoder: compute network weight using $L_{1}$ for $T=0$
end
      8) update discriminator and encoder weights
end
Algorithm 1 Pseudocode of DCAN
VI Conclusion
@@
We proposed DCAN, a non closed form approach to deep clustering over traditional closed form Euclidean or KL divergence loss function. The objective of traditional deep clustering is to optimize the autoencoder latent space with clustering information. The key to this technique lies in a loss function that minimizes both Kmeans partitioning and distribution of the samples in the latent space. Despite the recent success of deep clustering, there are some concerns: i) How to extend the inputs within the loss function to probabilistic models? ii) Can we backpropagate the neural network when there is no closed form learning for the probabilistic function. This paper addresses these two concerns both in theory and a proposed approach for a novel deep clustering approach based on adversarial net.
neural network objective cluster algo. method ACC
- Kmeans Kmeans 0.6815
recon. loss Kmeans AE 0.7373
recon. + Gaussian prior loss Kmeans VAE [10] 0.7663
recon. + clust. loss Kmeans ABC [7] 0.8872
Proposed Method
adversarial net based clust. loss Kmeans DCAN (deep features) 0.9047
Table II: SVHN
neural network objective cluster algo. method ACC
- Kmeans Kmeans [7] 0.674
- NCUT NCUT [7] 0.696
- Spectral Spectral [7] 0.693
recon. loss Kmeans AE[23] 0.6081
recon. + Euclidean dist. clust. loss Kmeans ABC[7] 0.715
KLD clust. loss Student’s t-dist. DEC [22] 0.6246
recon. + Euclidean dist. clust. loss Kmeans DC-Kmeans [21] 0.6442
recon. + Euclidean dist. clust. loss GMM DC-GMM [21] 0.6476
Proposed Method
adversarial net based clust. loss Kmeans DCAN (raw pixels) 0.8276
Table III: USPS
neural network objective cluster algo. method ACC
- Kmeans Kmeans [7] 0.535
- NCUT NCUT [7] 0.543
- Spectral Spectral [7] 0.556
recon. loss Kmeans AE [23] 0.6903
recon. + Euclidean dist. clust. loss Kmeans ABC [7] 0.760
recon. + Euclidean dist. clust. loss Kmeans DC-Kmeans [21] 0.8015
recon. + Euclidean dist. clust. loss GMM DC-GMM [21] 0.8555
KLD clust. loss Student’s t-dist. DEC [22] 0.843
recon. + Euclidean dist. clust. loss Kmeans DCN [14] 0.830
adversarial net based subspace loss Affinity matrix DASC [13] 0.804
self-augmented training RIM IMSAT-RPT [24] 0.896
pairwise-classification cosine distance DAC [25] 0.9775
ladder network ensemble KINGDRA-LADDER-1M [20] 0.985
Proposed Method
adversarial net based clust. loss Kmeans DCAN (raw pixels) 0.8565
adversarial net based clust. loss Kmeans DCAN (deep features) 0.995
Table IV: MNIST
neural network objective cluster algo. method ACC
KLD clust. loss Student’s t-dist. DEC [22] 0.469
self-augmented training RIM IMSAT-RPT [24] 0.455
pairwise-classification cosine distance DAC [25] 0.5218
ladder network ensemble KINGDRA-LADDER-1M [20] 0.546
Proposed Method
adversarial net based clust. loss Kmeans DCAN (deep features) 0.5844
Table V: CIFAR10
neural network objective clustering method ACC
- Kmeans Kmeans 0.6018
recon. + Euclidean dist. clust. loss Kmeans ABC 0.7019
recon. + Euclidean dist. clust. loss Kmeans DC-Kmeans [21] 0.7301
recon. + Euclidean dist. clust. loss GMM DC-GMM [21] 0.6906
KLD clust. loss Student’s t-dist. DEC [22] 0.7217
self-augmented training Kmeans IMSAT-RPT [24] 0.719
ladder network ensemble KINGDRA-LADDER-1M [20] 0.705
Proposed Method
adversarial net based clust. loss Kmeans DCAN (raw features) 0.7867
Table VI: Reuters-10k

### v3

@@ 2.1 Training DCAN
$\begin{array}[]{c}L_{DCAN}=\left\{\frac{1}{N}\sum_{n=1}^{N}\;\underset{x^{\left(n\right)}\sim p_{data}}{E}\left[\ln\left(1-D\left(G(x^{\left(n\right)})\right)\right)\right]\right\}_{T=0}\\ +\left\{\frac{1}{N}\sum_{n=1}^{N}\;\underset{z^{\left(n\right)}\sim p(z|\theta^{*})}{E}\left[\ln D\left(z^{\left(n\right)}\right)\right]\right\}_{T=1}\end{array}$ (1)
3 Experiments
@@
We selected some benchmarked datasets used by deep clustering for our experiments in Tables 2. There are several factors affecting deep clustering. i) The clustering algorithm used in the latent space e.g. Kmeans or GMM. ii) The number of hidden layers and nodes of the neural network. iii) The gradient ascent method used for training the loss functions. iv) The activation functions used. For DCAN, we use $tanh$ for the hidden layers and $sigmoid$ for the output. We set the clustering iteration to one in Table 2 and a sufficient statistics of at least 600 points in the raw latent space to estimate the cluster parameters. The number of cluster is the same as the number of classes. We use a minibatch size of $N=16$ samples, 500 iterations and we use gradient ascent with momentum with a learning rate between $10^{-3}$ and $10^{-5}$. We use accuracy (ACC) [15] to evaluate the performance of our clustering. MNIST: A well cited 10 digit classes dataset with no background but it has more samples than USPS. The train and test sample size is at 50K and 10K. Our settings are $196-384-256$ for the encoder and $256-16-1$ for the discriminator. The input dimension is originally at $28\times 28$ but we have downsampled it to $14\times 14$ and flattened to 196. In Table 2, DCAN using raw pixel information obtained $0.8565$ for ACC which is on par or better than deep clustering methods including DEC, ABC and DC-GMM. However on MNIST, most deep clustering methods could not obtain the performance of recent deep learning methods such as KINGDRA-LADDER-1M (a.k.a. KINGDRA) and IMSAT as their goals are much more ACC result oriented. For KINGDRA, it combines several ideas together including ladder network [16] and psuedo-label learning [17] for the neural network and using ensemble clustering. In DAC, their approach is based on pairwise classification. In IMSAT-RPT, the neural network was trained by augmenting the training dataset. In our case, we simplified our focused on deep clustering using adversarial net. DCAN penalizes samples for being an outlier from their cluster centers and does so uniquely by using the discriminator to penalize the encoder. The original objective of deep clustering is specifically aimed at clustering. In some datasets, the latent space of this objective do not necessarily co-serve as the most important factor for raw pixel feature extraction. Thus, we also used ResNet18 pretrained model as a feature extractor for MNIST, prior to deep clustering. This simple step allows the ACC for DCAN to improved to 0.995 (not shown in Table 2) and in fact outperforms KINGDRA. Reuters-10k: The Reuters-10k dataset according to [18, 19] contains 10K samples with 4 classes and the feature dimension is 2000. End-to-end learning is performed on this dataset. We compared DCAN to other unsupervised deep learning approaches. DCAN uses $2000-100-500$ and $500-100-1$ respectively for the encoder and discriminator. In Table 2, we observed that Kmeans alone can achieve an ACC of 0.6018 on the raw feature dimension. Most deep learning methods can achieve between 0.69 to 0.73 on this dataset. DCAN was able to obtain the best ACC at 0.7867. CIFAR10: A 10 classes object categorization dataset with 50K training and 10K testing of real images. On CIFAR10, most methods have difficulty performing end-to-end learning. As compared to deep ConvNet which utilizes convolutional neural network and supervised learning for a specific task (i.e. feature extraction), deep clustering typically rely on fewer hidden layer in the encoder for end-to-end learning (i.e. both feature extraction and clustering). Thus, we followed the experiment setup in KINGDRA [20] by using the “avg pool” layer of ResNet50 [21] (ImageNet pretrained) as a feature extractor with a dimension of 2048. In Table 2, DCAN uses $2048-1024-512-128$ and $128-16-1$ respectively for the encoder and discriminator. Despite the stronger performance of DAC, IMSAT and KINGDRA on MNIST, we outperformed these comparisons on CIFAR10 at $0.5844$. This is despite the fact that both KINGDRA and IMSAT also use ResNet50 (ImageNet pretrained) for CIFAR10.
Approach MNIST Reuters10k CIFAR10
ABC [7] 0.760 0.7019 0.435
DEC [19] 0.843 0.7217 0.469
DC-GMM [18] 0.8555 0.6906 -
AAE [13] 0.8348 0.6982 -
IMSAT-RPT [22] 0.896 0.719 0.455
KINGDRA [20] 0.985 0.705 0.546
DCAN (proposed) 0.8565 0.7867 0.5844
Table 2: ACC Benchmark
4 Conclusion
@@
The objective of deep clustering is to optimize the autoencoder latent space with clustering information. The key to this technique lies in a loss function that minimizes both clustering and encoder in the latent space. Despite the recent success of deep clustering, there are some concerns: i) How to use probabilistic approach for the loss function? ii) Can we backpropagate the neural network when there is no closed form solution for the probabilistic function? A JS divergence approach does not suffer from asymmetry like KL divergence loss. Unlike the latter, there is no closed form solution to the JS divergence when backpropagating the network. This paper addresses the above two concerns by proposing a novel deep clustering approach based on adversarial net to estimate JS divergence for deep clustering.